# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### 08 Dec 2023 v1

@@ 1 Introduction
Large language models (LLMs) has amazed the world with their astonishing abilities and performance in solving a wide range of problems [4, 29, 7, 56, 46, 47]. This is accompanied by excessive costs and carbon emissions for training and deploying these models, as their sizes grow rapidly in recent years. In general, costs for inference are dominant in the long run, as each model will be deployed to solve many problems for a long period of time. This has inspired researchers and engineers to develop various approaches for accelerating LLM inference.
The focus of this work is early exiting, which accelerates inference by allowing a deep neural network to make predictions and exit early in the network for certain inputs, without running the forward pass through the full network. This is achieved by augmenting a standard neural network architecture (with a single exit at the end) with additional early-exit layers that transform intermediate hidden states into early outputs. The early-exit model architecture, as visualized in Figure 1, not only retains the full capacity of a large model, but is also capable of adaptively using a smaller amount of computation for solving simpler problems. The idea of early exiting is a natural analogue of how human speaks, thinks, and make decisions: not every problem requires or deserves the same amount of consideration, and one shall opt for fast reaction to simple problems without overthinking. Early exiting has been an active research area and widely applied in natural language processing, computer vision, and other areas [13, 24, 15, 59, 11, 38, 51, 37, 52, 45, 16, 18, 14, 53]. More recently, it starts to gain attention in the LLM domain [37, 9, 2, 48], and is recognized as a promising direction for further reducing the latency and costs of LLM inference [31].
@@ 1.1 Goal and motivations
The primary goal of this work is to build the infrastructure for scaling up training and inference of early-exit LLMs. This is motivated by the observation that the sizes of early-exit models in prior works are still relatively small. The largest early-exit LLM that we are aware of is an OPT model [56] with 6.7 billion (6.7B) parameters [9]; in contrast, standard LLMs at much larger scales, e.g. the 175B GPT-3 [4], 530B Megatron-Turing NLG [42], 540B PaLM [7], or even larger sparsely activated models, have been well trained and deployed in many applications. It is an urgent need for the community to truly understand the efficacy of early exiting for LLMs at larger scales, which is indispensable towards making early exiting a useful and practical option in complex scenarios that only sufficiently large LLMs can handle.
@@ 1.2 Challenges
Moreover, training efficiency for early-exit LLMs requires specific design. While sizes of early-exit layers are often regarded as negligible for many neural network architectures, this is not the case for LLMs, where each early exit contains (at least) a large output embedding matrix that transforms hidden states into logits on the vocabulary. A naive implementation of early-exit LLM training can cause large computational overhead compared to standard LLM training.
@@ 1.3 Main contributions
We propose EE-LLM, a system for large-scale training and inference of early-exit (EE) LLMs with 3D parallelism, which is designed to tackle the aforementioned challenges. EE-LLM is built upon Megatron-LM [41, 28, 20, 42], and augments it with various functionalities for early exiting. In addition to compatibility with existing functionalities of 3D parallelism provided by Megatron-LM, EE-LLM also implements a variety of algorithmic innovations, including a lightweight method that facilitates backpropagation for the early-exit training objective through pipeline stages, various techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching (with one based on a novel form of pipeline parallelism and another based on KV recomputation). Implementation of EE-LLM has been well optimized for maximum training and inference efficiency. Our analytical and empirical study confirms that, with negligible computational overhead (caused by early-exit layers) during training with 3D parallelism, one obtains an early-exit LLM that generates tokens with adaptive token-wise exit selection, achieving outstanding speedup without compromising output quality during inference. The source code for EE-LLM can be found at https://github.com/pan-x-c/EE-LLM.
@@ 2.1 Transformers
The Transformer architecture [49] has been playing a dominant role in natural language processing (NLP) and large language models (LLMs) [3, 57]. It is typically composed of an input embedding layer, a stack of Transformer layers, and finally an output layer. Each Transformer layer consists of cross-attention and/or self-attention modules, a multi-layer perceptron (MLP), and layer normalization (LayerNorm [1] or RMSNorm [55]). Transformers can be categorized into three types: encoder-only, encoder-decoder, and decoder-only. For the latter two, there is an output embedding matrix in the output layer, which transforms hidden states into logits on a (typically large) vocabulary that can be used for generating tokens. An LLM can be learned by unsupervised pre-training, e.g. minimizing the negative log-likelihood of next-token prediction on a large corpus [33, 34]. In this work, we focus on the decoder-only generative pre-training (GPT) Transformer architecture [33, 34], though many of our ideas are widely applicable to other Transformer architectures or generic deep neural networks.
@@ 2.2 Early-exit LLMs
The standard way of training an early-exit model is to minimize a weighted sum of early-exit and final-exit training losses [38, 36]. Note that early-exit layers bring additional computational overhead to training. This is especially the case for LLMs, primarily due to the large output embedding matrix of size $h\times V$ within each early-exit layer, where $h$ is the hidden dimension and $V$ is the vocabulary size. We call an early-exit layer minimalistic if it has the same structure as the final output layer of the GPT model architecture [33, 34], i.e. an output embedding matrix, plus an optional LayerNorm/RMSNorm in front of it. Additional modules can be added to early-exit layers for increased expressivity and adaptivity of early exits.
@@ 2.3 Megatron-LM and 3D parallelism
Pipeline parallelism [26, 27, 12, 21], on the other hand, partitions a deep model along the depth dimension into multiple pipeline stages. Moreover, each data batch is divided into multiple microbatches, and their forward/backward computational tasks are scheduled among those multiple pipeline stages. More specifically, each stage performs the forward computation for each microbatch and sends the resulted hidden states to another stage; later on, it performs the backward computation for the same microbatch after receiving the gradients (of the training objective) with respect to the sent hidden states. Pipeline parallelism only requires sparse and inexpensive point-to-point (P2P) communication between pipeline stages, which makes it applicable and oftentimes must-have in much broader scenarios when tensor (and sequence) parallelism is infeasible or insufficient, whether in GPU clusters or in decentralized settings [54]. The main concern with pipeline parallelism is its low utilization rate of computational resources, due to pipeline bubbles and load imbalance across pipeline stages [28]; this will be discussed in more details in Section 4.2.
@@ 3.2 Training
EE-LLM contains the essential functionalities for training early-exit LLMs, which tackle the main challenges outlined in Section 1.2, i.e. how to train with 3D parallelism (especially pipeline parallelism) while minimizing the computational overhead (compared to standard LLM training) caused by early-exit layers. In addition to engineering efforts for compatibility with existing functionalities in Megatron-LM, we design and implement a simple yet effective algorithm that facilitates pipeline parallelism with multiple (early-exit or final-output) training losses located on different pipeline stages, which is not possible in standard pipeline parallelism implemented by Megatron-LM or other frameworks. Moreover, our analytical and empirical study shows that training an early-exit LLM with EE-LLM is almost as efficient as training a standard LLM, in terms of training time and peak GPU memory; this is achieved by various performance optimizations that we design and implement in EE-LLM, especially the ones that (passively or actively) leverage idle computational resources in pipeline parallelism. Finally, EE-LLM contains some advanced features for more fine-grained control or optimization of the training procedure, including the option of changing early-exit loss weights during training, and a novel method of further improving resource utilization by filling pipeline bubbles with useful computation.
@@ 4 Training
In this section, we first introduce (in Section 4.1) the essentials of scaling up early-exit LLM training with massive 3D parallelism; in particular, we demonstrate how to leverage pipeline parallelism via a novel approach of executing backpropagation for the early-exit training loss through pipeline stages. Next, we show (in Section 4.2) how to maximize training efficiency via various performance optimizations, e.g. choosing wisely where to add early exits, and scheduling computation in the correct order. By leveraging idle computational resources in pipeline parallelism, training an early-exit LLM bears negligible overhead compared to training a standard LLM, as confirmed by our analytical and empirical study. Finally, we explore (in Section 4.3) some advanced features implemented in EE-LLM, including the option to change the weights of early-exit losses during the training process, and a novel approach of improving hardware utilization by filling pipeline bubbles with useful computation for additional microbatches.
@@ 4.1 Backpropagation through pipeline stages
Our implementation of training, i.e. optimizing the loss function in Eq. (1), is compatible with all types of parallelism in Megatron-LM. Indeed, with some engineering efforts, existing functionalities in Megatron-LM for data and tensor/sequence parallelism are directly applicable. The major challenge lies in pipeline parallelism, since it is not immediately clear how to calculate gradients for Eq. (1) via backpropagation through pipeline stages. In a single-GPU scenario with vanilla PyTorch, one simply needs to define loss as the weighted sum of all early-exit and final-exit losses, and then run loss.backward() for gradient calculation. This is obviously not feasible with pipeline parallelism, since losses $\{{\mathcal{L}}_{i}^{\mathsf{exit}}\}$ are located on different pipeline stages, and there is only limited point-to-point (P2P) communication between each pair of adjacent stages (as explained in Section 2.3). On the other hand, Megatron-LM only supports backpropagation for a single loss function defined in the last pipeline stage.
@@ 4.1.1 Methodology
To tackle this challenge, we propose a simple yet effective algorithm that instructs each pipeline stage to calculate the desired gradients correctly, without any additional communication overhead between stages. Let us first re-write the early-exit training loss defined in Eq. (1) as
@@
Figure 2: A high-level visualization for the process of backpropagation (including forward and backward passes) for an early-exit model partitioned into four pipeline stages.
@@ 4.2 Training efficiency
In the following, we analyze the training time and peak GPU memory of training an early-exit model with pipeline parallelism, and propose some performance optimizations. We refer readers to the literature (e.g. the Megatron-LM series [41, 28, 20]) for thorough study of training efficiency with 3D parallelism; we will use that as a baseline, and focus on the additional computational overhead caused by early exits.
Let us first identify the major sources of low resource utilization in the original 1F1B pipeline schedule for training a standard LLM (visualized in Figure 3(a)), which lays the foundation for our analysis later.
@@
  • Implicit bubbles, i.e. dark gray areas in Figure 3(a). This is caused by load imbalance across pipeline stages (even though Transformer layers are evenly divided into stages in Megatron-LM) [^3] For simplicity, we do not consider other flexible ways of dividing a LLM into pipeline stages, as dividing Transformer layers evenly remains one of the most practical and robust option in practice. We also do not consider the case where the input/output embedding matrix occupies a separate pipeline stage, although our techniques for training an early-exit LLM can be generalized to this case in principle.: the first stage has the additional computation for the input embedding layer, and (more importantly) the last pipeline stage has the additional computation for the output logits (via the output embedding layer) as well as the training loss. For LLMs, these additional computational costs are not negligible, primarily due to large vocabulary sizes.
  • Idle memory. Memory usage is also imbalanced across pipeline stages: earlier stages in the 1F1B schedule have to save the intermediate activations for more microbatches, and the first/last stage has to save an extra embedding matrix. As a result, the first stage is typically the bottleneck of peak memory, while later stages have idle memory [20, Section 4.2.3].
Figure 3: One iteration of the 1F1B pipeline schedule in different cases. Within one iteration, each stage goes through a warm-up phase (forward steps of the beginning microbatches), a steady 1F1B (one-forward-one-backward) phase, and a cool-down phase (backward steps of the final microbatches). “Backbone forward/backward” stands for computation of Transformer layers on the backbone, while “Exit forward/backward” stands for computation of early-exit or final-exit layers. The number in each block denotes the index of the corresponding microbatch. Critical paths are marked by dashed red lines. In this figure, the degree of pipeline parallelism is set to 4, and each data batch is divided into 6 microbatches. It is assumed that for each stage, the ratio of time between forward computation for the backbone and forward computation for the (early/final-)exit layer is 2:1; moreover, we adopt the common assumption that the ratio of time between forward and backward computation is 1:2. Note that our analytical study does not rely on the specific values of these ratios. For clarify, we ignore computation related to the input embedding layer, as well as P2P communication latency between pipeline stages.
@@ 4.2.1 Utilization of idle resources
More specifically, the overhead to training time caused by additional early exits can be negligible. If we choose $k$ middle stages and add one minimalistic early-exit layer to each of them, then the training time per iteration, i.e. time for processing one data batch, will (in theory) increase only by $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$, where $f_{\mathsf{EE}}$ and $b_{\mathsf{EE}}$ represent the time needed for one forward and backward pass of one microbatch for one minimalistic early-exit layer, respectively. To see this, first notice that the computation of early-exit layers in the steady 1F1B phase can be perfectly fit into the implicit bubbles. Therefore, the critical path remains the same as in the case without early exits, which consists of (1) the forward steps on all stages for the first microbatch, (2) the 1F1B steady phase on the last stage, and (3) the backward steps on all stages for the last microbatch. The early-exit layers, located separately on $k$ middle stages, will cause a $k\times f_{\mathsf{EE}}$ overhead to the first part of the critical path, and a $k\times b_{\mathsf{EE}}$ overhead to the third part, leading to the aforementioned claim on the overhead to training time. See Figure 3(a) and (b) for a visualization.
4.2.2 Further performance optimization

### 01 Feb 2024 v2

@@ 1 Introduction
Large language models (LLMs) have amazed the world with their astonishing abilities and performance in solving a wide range of problems [8, 50, 11, 89, 74, 75]. This is accompanied by excessive costs and carbon emissions for training and deploying these models, as their sizes grow rapidly in recent years. In general, costs for inference are dominant in the long run, as each model will be deployed to solve many problems for a long period of time. This has inspired researchers and engineers to develop various approaches for accelerating LLM inference.
The focus of this work is early exiting, which accelerates inference by allowing a deep neural network to make predictions and exit early in the network for certain inputs, without running the forward pass through the full network. This is achieved by augmenting a standard neural network architecture (with a single exit at the end) with additional early-exit layers that transform intermediate hidden states into early outputs. The early-exit model architecture, as visualized in Figure 1, not only retains the full capacity of a large model, but is also capable of adaptively using a smaller amount of computation for solving simpler problems. The idea of early exiting is a natural analogue of how human speaks, thinks, and make decisions: not every problem requires or deserves the same amount of consideration, and one shall opt for fast reaction to simple problems without overthinking [31]. Early exiting has been an active research area and widely applied in natural language processing, computer vision, and other areas [23, 45, 25, 94, 19, 63, 81, 62, 82, 42, 26, 73, 27, 31, 38, 60, 24, 83, 14]. More recently, it starts to gain attention in the LLM domain [62, 13, 3, 76], and is recognized as a promising direction for further reducing the latency and costs of LLM inference [53].
@@ 1.1 Goal and motivations
The primary goal of this work is to build the infrastructure for scaling up training and inference of early-exit LLMs. This is motivated by the observation that the sizes of early-exit models in prior works are still relatively small. The largest early-exit LLM that we are aware of is a Llama 2 model [75] with 13 billion (13B) parameters [76]; in contrast, standard LLMs at much larger scales, e.g. the 175B GPT-3 [8], 530B Megatron-Turing NLG [68], 540B PaLM [11], or even larger sparsely activated models, have been well trained and deployed in many applications. It is an urgent need for the community to truly understand the efficacy of early exiting for LLMs at larger scales, which is indispensable for making early exiting a useful and practical option in complex scenarios that only sufficiently large LLMs can handle.
@@ 1.2 Challenges
Moreover, training efficiency for early-exit LLMs requires special design. While sizes of early-exit layers are often regarded as negligible for many neural network architectures, this is not the case for LLMs, where each early exit contains (at least) a large output embedding matrix that transforms hidden states into logits on the vocabulary. A naive implementation of early-exit LLM training can cause large computational overhead compared to standard LLM training.
@@ 1.3 Main contributions
We propose EE-LLM, a system for large-scale training and inference of early-exit (EE) LLMs with 3D parallelism, which is designed to tackle the aforementioned challenges. EE-LLM is built upon Megatron-LM [67, 49, 36, 68], and augments it with various functionalities for early exiting. In addition to compatibility with existing functionalities of 3D parallelism provided by Megatron-LM, EE-LLM also implements a variety of algorithmic innovations, including a lightweight method that facilitates backpropagation for the early-exit training objective through pipeline stages, various techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching (with one based on a novel form of pipeline parallelism and another based on KV recomputation). Implementation of EE-LLM has been well optimized for maximum training and inference efficiency. Our analytical and empirical study confirms that, with negligible computational overhead (caused by early-exit layers) during training with 3D parallelism, one obtains an early-exit LLM that generates tokens with adaptive token-wise exit selection, achieving outstanding speedup without compromising output quality during inference. In other words, EE-LLM facilitates training and inference of early-exit LLMs that are as large as the maximum sizes of standard LLMs allowed by Megatron-LM, given the same amount of computational resources. The source code for EE-LLM can be found at https://github.com/pan-x-c/EE-LLM.
@@ 2.1 Transformers
The Transformer architecture [77, 71] has been playing a dominant role in natural language processing (NLP) and large language models (LLMs) [7, 90]. It is typically composed of an input embedding layer, a stack of Transformer layers, and finally an output layer. Each Transformer layer consists of cross-attention and/or self-attention modules [4, 34, 52], a multi-layer perceptron (MLP), and layer normalization (LayerNorm [2] or RMSNorm [88]). Transformers can be categorized into three types: encoder-only, encoder-decoder, and decoder-only. For the latter two, there is an output embedding matrix in the output layer, which transforms hidden states into logits on a (typically large) vocabulary that can be used for generating tokens. An LLM can be learned by unsupervised pre-training, e.g. minimizing the negative log-likelihood of next-token prediction on a large corpus [55, 56]. In this work, we focus on the decoder-only generative pre-training (GPT) Transformer architecture [55, 56], though many of our ideas are widely applicable to other Transformer architectures or generic deep neural networks.
@@ 2.2 Early-exit LLMs
The standard way of training an early-exit model is to minimize a weighted sum of early-exit and final-exit training losses [63, 61]. Note that early-exit layers bring additional computational overhead to training. This is especially the case for LLMs, primarily due to the large output embedding matrix of size $h\times V$ within each early-exit layer, where $h$ is the hidden dimension and $V$ is the vocabulary size. We call an early-exit layer minimalistic if it has the same structure as the final output layer of the GPT model architecture [55, 56], which includes an output embedding matrix, plus an optional LayerNorm/RMSNorm in front of it. Additional modules can be added to early-exit layers for increased expressivity and adaptivity of early exits.
@@ 2.3 Megatron-LM and 3D parallelism
Pipeline parallelism [47, 48, 20, 41], on the other hand, partitions a deep model along the depth dimension into multiple pipeline stages. Moreover, each data batch is divided into multiple microbatches, and their forward/backward computational tasks are scheduled among those multiple pipeline stages. More specifically, each stage performs the forward computation for each microbatch and sends the resulted hidden states to another stage; later on, it performs the backward computation for the same microbatch after receiving the gradients of the training objective with respect to the sent hidden states. Pipeline parallelism only requires sparse and inexpensive point-to-point (P2P) communication between pipeline stages, which makes it applicable and oftentimes must-have in much broader scenarios when tensor (and sequence) parallelism is infeasible or insufficient, whether in GPU clusters or in decentralized settings [86, 84]. The main concern with pipeline parallelism is its low utilization rate of computational resources, due to pipeline bubbles and load imbalance across pipeline stages [49]; this will be discussed in more details in Section 4.2.
@@ 3.2 Training
EE-LLM contains the essential functionalities for training early-exit LLMs, which tackle the main challenges outlined in Section 1.2, i.e. how to train with 3D parallelism (especially pipeline parallelism) while minimizing the computational overhead (compared to standard LLM training) caused by early-exit layers. In addition to substantial engineering efforts for compatibility with existing functionalities in Megatron-LM, we design and implement a simple yet effective algorithm that facilitates pipeline parallelism with multiple early-exit or final-output training losses located on different pipeline stages, which is not possible in standard pipeline parallelism implemented by Megatron-LM or other frameworks. Moreover, our analytical and empirical study shows that training an early-exit LLM with EE-LLM is almost as efficient as training a standard LLM, in terms of training time and peak GPU memory; this is achieved by various performance optimizations that we design and implement in EE-LLM, especially the ones that passively or actively leverage idle computational resources in pipeline parallelism. Finally, EE-LLM contains some advanced features for more fine-grained control or optimization of the training procedure, including the option of changing early-exit loss weights during training, and a novel method of further improving resource utilization by filling pipeline bubbles with useful computation.
@@ 4 Training
In this section, we first introduce (in Section 4.1) the essentials of scaling up early-exit LLM training with massive 3D parallelism. In particular, we demonstrate how to leverage pipeline parallelism via a novel approach of executing backpropagation for the early-exit training loss through pipeline stages. Next, we show (in Section 4.2) how to maximize training efficiency via various performance optimizations, e.g. choosing wisely where to add early exits, and scheduling computation in the correct order. By leveraging idle computational resources in pipeline parallelism, training an early-exit LLM bears negligible overhead compared to training a standard LLM, as confirmed by our analytical and empirical study. Finally, we explore (in Section 4.3) some advanced features implemented in EE-LLM, including the option to change the weights of early-exit losses during the training process, and a novel approach of improving hardware utilization by filling pipeline bubbles with useful computation for additional microbatches.
@@ 4.1 Backpropagation through pipeline stages
Our implementation of training, i.e. optimizing the loss function in Eq. (1), is compatible with all types of parallelism in Megatron-LM. Indeed, with some engineering efforts, existing functionalities in Megatron-LM for data and tensor/sequence parallelism are directly applicable. The major challenge lies in pipeline parallelism, since it is not immediately clear how to calculate gradients for Eq. (1) via backpropagation through pipeline stages. In a single-GPU scenario with vanilla PyTorch, one simply needs to define loss as the weighted sum of all early-exit and final-exit losses, and then run loss.backward() for gradient calculation. This is obviously not feasible with pipeline parallelism, since losses $\{{\mathcal{L}}_{i}^{\mathsf{exit}}\}$ are located on different pipeline stages, and there is only limited P2P communication between each pair of adjacent stages (as explained in Section 2.3). On the other hand, Megatron-LM only supports backpropagation for a single loss function defined in the last pipeline stage.
@@ 4.1.1 Methodology
To tackle this challenge, we propose a simple yet effective algorithm that instructs each pipeline stage to calculate the desired gradients correctly, without any additional communication overhead between stages. To explain our method, let us first re-write the early-exit training loss defined in Eq. (1) as
@@
Figure 2: A high-level visualization for the process of backpropagation, including forward and backward passes, for an early-exit model partitioned into four pipeline stages.
@@ 4.2 Training efficiency
In the following, we analyze the training time and peak GPU memory of training an early-exit model with pipeline parallelism, and propose some performance optimizations. We refer readers to the literature (e.g. the Megatron-LM series [67, 49, 36]) for thorough study of training efficiency with 3D parallelism for standard LLMs without early exits; we will use that as a baseline, and focus on the additional computational overhead caused by early exits.
Let us first identify the major sources of low resource utilization in the original 1F1B pipeline schedule for training a standard LLM, which lays the foundation for our analysis later.
@@
  • Implicit bubbles, i.e. dark gray areas in Figure 3(a). This is caused by load imbalance across pipeline stages, even though Transformer layers are evenly divided into stages in Megatron-LM [^3] For simplicity, we do not consider other flexible ways of dividing a LLM into pipeline stages, as dividing Transformer layers evenly remains one of the most practical and robust option in practice. We also do not consider the case where the input/output embedding matrix occupies a separate pipeline stage, although our techniques for training an early-exit LLM can be generalized to this case in principle.. In particular, the first stage has the additional computation for the input embedding layer, and more importantly, the last pipeline stage has the additional computation for the output logits (via the output embedding layer) as well as the training loss. For LLMs, these additional computational costs are not negligible, primarily due to large vocabulary sizes.
  • Idle memory. Memory usage is also imbalanced across pipeline stages: earlier stages in the 1F1B schedule have to save the intermediate activations for more microbatches, and the first/last stage has to save an extra input/output embedding layer, plus the corresponding gradients and optimizer states. As a result, the first stage is typically the bottleneck of peak memory usage, while later stages have idle memory [36, Section 4.2.3].
Figure 3: One iteration of the 1F1B pipeline schedule in different cases. Within one iteration, each stage goes through a warm-up phase (forward steps of the beginning microbatches), a steady 1F1B (one-forward-one-backward) phase, and a cool-down phase (backward steps of the final microbatches). “Backbone forward/backward” stands for computation of Transformer layers on the backbone, while “Exit forward/backward” stands for computation of early-exit or final-exit layers. The number in each block denotes the index of the corresponding microbatch. Critical paths are marked by dashed red lines. In this figure, the degree of pipeline parallelism is set to 4, and each data batch is divided into 6 microbatches. It is assumed that for each stage, the ratio of time between forward computation for the backbone and forward computation for the early/final-exit layer is 2:1; moreover, we adopt the common assumption that the ratio of time between forward and backward computation is 1:2. Note that our analytical study does not rely on the specific values of these ratios. For clarity, we ignore computation related to the input embedding layer, as well as P2P communication latency between pipeline stages.
@@ 4.2.1 Utilization of idle resources
More specifically, the overhead to training time caused by additional early exits can be negligible. If we choose $k$ middle stages and add one minimalistic early-exit layer to each of them, then the training time per iteration, i.e. time for processing one data batch, will (in theory) increase only by $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$, where $f_{\mathsf{EE}}$ and $b_{\mathsf{EE}}$ represent the time needed for one forward and backward pass of one microbatch for one minimalistic early-exit layer, respectively. To see this, first notice that the computation of early-exit layers in the steady 1F1B phase can be perfectly fit into the implicit bubbles. Therefore, the critical path remains the same as in the case without early exits, which consists of (1) the forward steps on all stages for the first microbatch, (2) the 1F1B steady phase on the last stage, and (3) the backward steps on all stages for the last microbatch. The early-exit layers, located separately on $k$ middle stages, will cause a $k\times f_{\mathsf{EE}}$ overhead to the first part of the critical path, and a $k\times b_{\mathsf{EE}}$ overhead to the third part, leading to the aforementioned claim on the overhead to training time. See Figures 3(a) and (b) for a visualization.
4.2.2 Further performance optimizations